Circular No. 5757

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

V1974 CYGNI H. Dinerstein and R. Benjamin communicate: “Infrared spectroscopy

of V1974 Cyg (N Cyg 1992) by D. Lester, N. Gaffney, and B. Smith at the McDonald Observatory 2.7-m reflector shows the appearance of new infrared coronal lines, [Al IX] 2.040 μm and [S IX] 1.250 μm. The ionization energy to create these ions are 285 and 328 eV, respectively. Both lines are clearly resolved from neighboring lines. The [S IX] 1.250-μm line was tentatively detected by Woodward and Greenhouse (IAUC 5653) with a reported line strength of 0.08 times that of the adjacent Paβ line. This strength was consistent, however, with an identification as He I 1.253 μm, which had been observed as early as 1992 June, before the onset of the coronal phase. On 1993 Apr. 2.4 UT, we find the strength of the 1.25-μm feature is a factor of 8 too strong to be due to He I alone; it has increased to ∼ 0.6 times the strength of Paβ. Therefore, we regard it as highly likely that [S IX] is now present. Further confirmation that the level of ionization in the ejecta has continued to increase since our previous epoch of observations (1992 Dec. 10) is seen in the ratio [Si VII] 2.47-μm/[Si VI] 1.96-μm, which increased from ∼ 1.6 to 3.5 between 1992 December and 1993 April, while the absolute fluxes of both lines dropped. These new infrared spectra supplement and confirm the recent report of very high ionization species ([Fe X], [Ar X], and possibly [Fe XIV]) in the optical spectrum of V1974 Cyg by Garnavich (IAUC 5746).”

SUPERNOVA 1993J IN NGC 3031 G. F. Lawrence and A. Paulson report the following infrared photom-

etry of SN 1993J obtained on Apr. 3.39 UT at the University of Minnesota 0.76-m telescope using an InSb photometer (standard infrared photometric techniques were employed with a 27′′ beamsize and a 44′′ north/south throw; calibration was against the IRTF standard stars HR 3888 and HR 5447): J = 10.91 ± 0.05, H = 10.66 ± 0.05, K = 10.48 ± 0.08.

Visual magnitude estimates: Mar. 10.27 UT, [15.5 (W. Wren, McDonald Observatory); Apr. 2.17, 11.4 (C. Spratt, Victoria, BC); 3.28, 11.5 (W. Johnson, Anza, CA); 4.03, 11.6 (P. Schmeer, Bischmisheim, Germany); 5.03, 11.8 (J. Bortle, Stormville, NY); 5.271, 11.8 (R. Royer, Lakewood, CA); 6.26, 11.7 (W. G. Dillon, Missouri City, TX); 7.23, 11.6 (Wren); 8.08, 11.4 (C. Heil, Andover, MA); 9.20, 11.5 (Dillon); 9.25, 11.4 (Royer).

1993 April 9 Daniel W. E. Green